# The SPHEREx Satellite Mission

### v1

@@ 4.6 Data Processing and Archiving
The SPHEREx data processing flow includes four levels of data processing. Level 1 to 3 are performed by the SPHEREx Science Data Center (SSDC) at IPAC and Level 4 is performed by the SPHEREx Science Team. Importantly, the SSDC pipeline does not perform blind source detection and extraction. Instead, it measures the flux of sources listed in a reference catalog (Yang et al., 2025) that is generated by the SPHEREx Science Team which includes calibration sources (both photometric and astrometric), science targets for the cosmology and interstellar ices themes, and any nearby sources that need to be included for photometric separation. Figure 14 shows the high-level organization of the pipeline levels, data products and tools. Additional details for the Level 1 to 3 pipeline can be found in Akeson et al. (2025a) and the SPHEREx Explanatory Supplement (Akeson et al., 2025b).
@@ 4.6.2 Level 2 Processing
Correcting for detector non-linearity, which arises from declining gain as charge accumulates within each H2RG pixel (Zengilowski et al., 2020).
@@
In addition, the Level 2 pipeline derives the absolute gain matrix by combining both the flat-field and the dark current matrices with measured calibration factors. As described in §3.4.3, we derive the flat-field and dark current matrices using hundreds of input images obtained over varying Zodiacal light levels. We use observations of pre-selected primary calibrator stars to derive the calibration factors by comparing the measured and model flux as a function of wavelength (see Akeson et al. (2025b) for the list). We apply the calibration per pixel using the absolute gain matrix and the dark current calibration products. The pipeline determines the sub-pixel PSF in a grid over each band by stacking and deconvolving the measured response for a large number of isolated and high-SNR stars. The resulting product gives an effective PSF in each region that includes both the intrinsic optical response and the average blurring due to pointing jitter. The original design separated the optical PSF and the per-observation jitter, but the measured jitter is so low (see Figure 16) that we only use the average PSF per region. IRSA provides the absolute gain matrix, the dark current matrix and the derived PSFs as calibration products. The PSFs are also available as an extension within each Level 2 calibrated file.
@@ 5.2 Thermal Performance
Once the decontamination heaters were turned off, the instrument radiatively cooled to operating temperature as shown in Figure 15. The multi-stage thermal design achieved the expected temperatures of 290 K, 150 K, and 90 K on the 3 photon shields. Once the MWIR focal plane reached 45 K, we engaged a control heater to stabilize its temperature. The telescope cooled more slowly, as expected. Once it reached 62 K, we engaged a second control heater to stabilize the SWIR focal plane temperature. After cooling stabilized, we observed that the control heaters maintained these temperatures with 47 and 234 mW dissipation for the MWIR and SWIR focal planes, respectively.
@@ 5.5 Gain and Dark Current
We produced preliminary in-flight estimates of the flat-field gain and dark current using an ensemble of images taken over varying Zodiacal brightness levels (see §3.4.3). Both estimates closely resemble measurements obtained during laboratory testing of the instrument and focal planes. As we refine the algorithms and include more data, Akeson et al. (2025b) will contain the most recent estimates.
@@ 5.8 Stray Light Response
We are in the process of quantifying a range of stray light phenomena from in-flight observations (Dowell et al., 2026). Although there are a few exceptions, the effects and levels are generally in accordance with a prelaunch design study of the stray light properties. Within the optical field of view, we can readily characterize the extended PSF to enable PSF subtraction and source masking. We observe scattered light from bright stars in narrow regions close to the frame edges, and ghost images caused by reflections in the beam splitter that are most noticeable near the 2.4 m transition.
@@ 5.9.1 Atmospheric Emission
We observe bright He line emission (Kulkarni, 2025; Brammer et al., 2014) at 1.083 m in both bands 1 and 2. The amplitude varies over the orbit and with time, ranging from a few to 50 times brighter than the Zodiacal light level. Photon noise from the line unfortunately reduces sensitivity in this band (see Figure 19). We detect low levels of spectral leakage through the LVFs from the He line, most evident at 0.84, 1.25 and 1.49 m. Other atmospheric lines are much fainter than He but detectable, most notably the atomic oxygen lines at 0.78, 0.84 and 1.13 m.
@@ 5.9.2 Shuttle Glow
Shuttle glow arises from the spacecraft interacting with the upper atmosphere. Atomic oxygen striking the leading surfaces in the direction of travel produces a glow in the optical and infrared (Viereck et al., 1991) due to excited (Murad, 1998). Emission from other molecules have been proposed in the literature, including and , which may depend on the composition of the materials on the spacecraft (Holtzclaw et al., 1994) that chemically react with atomic oxygen. As noted in §4.2.1, we observed that shuttle glow increased strongly when the observatory was tipped forward along the direction of orbital motion. By restricting the range of zenith angles in the ram direction, we have reduced shuttle glow so that it is generally fainter than the Zodiacal sky, though a residual brightness remains.
@@ 6 Conclusions
Part of the research was carried out at the Jet Propulsion Laboratory and the California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The project thanks the many people who helped develop and implement hardware for the mission, including but not limited to staff at NASA, JPL, BAE, IPAC, IRSA, KASI, SpaceX, VSFB, Caltech, Viavi, Applied Aerospace Structures, UC Berkeley SSL, and Teledyne.

### v2

@@ 4.6 Data Processing and Archiving
The SPHEREx data processing flow includes four levels of data processing. Level 1 to 3 are performed by the SPHEREx Science Data Center (SSDC) at IPAC and Level 4 is performed by the SPHEREx Science Team. Importantly, the SSDC pipeline does not perform blind source detection and extraction. Instead, it measures the flux of sources listed in a reference catalog (Yang2025) that is generated by the SPHEREx Science Team which includes calibration sources (both photometric and astrometric), science targets for the cosmology and interstellar ices themes, and any nearby sources that need to be included for photometric separation. Figure 14 shows the high-level organization of the pipeline levels, data products and tools. Additional details for the Level 1 to 3 pipeline can be found in Akeson25b and the SPHEREx Explanatory Supplement (akeson25).
@@ 4.6.2 Level 2 Processing
Correcting for detector non-linearity, which arises from declining gain as charge accumulates within each H2RG pixel (Zengilowski20).
@@
Providing an estimate of the diffuse Zodiacal light (Crill2025) based on the spatial location of the image. The empirical model is based on Kelsall98, using spectral measurements from Tsumura10; Tsumura13.
@@
In addition, the Level 2 pipeline derives the absolute gain matrix by combining both the flat-field and the dark current matrices with measured calibration factors. As described in §3.4.3, we derive the flat-field and dark current matrices using hundreds of input images obtained over varying Zodiacal light levels. We use observations of pre-selected primary calibrator stars to derive the calibration factors by comparing the measured and model flux as a function of wavelength (see akeson25 for the list). We apply the calibration per pixel using the absolute gain matrix and the dark current calibration products. The pipeline determines the sub-pixel PSF in a grid over each band by stacking and deconvolving the measured response for a large number of isolated and high-SNR stars. The resulting product gives an effective PSF in each region that includes both the intrinsic optical response and the average blurring due to pointing jitter. The original design separated the optical PSF and the per-observation jitter, but the measured jitter is so low (see Figure 16) that we only use the average PSF per region. IRSA provides the absolute gain matrix, the dark current matrix and the derived PSFs as calibration products. The PSFs are also available as an extension within each Level 2 calibrated file.
@@ 5.2 Thermal Performance
Once the decontamination heaters were turned off, the instrument radiatively cooled to operating temperature as shown in Figure 15. The multi-stage thermal design achieved the expected temperatures of 290 K, 150 K, and 90 K on the 3 photon shields. Once the MWIR focal plane reached 45 K, we engaged a control heater to stabilize its temperature. The telescope cooled more slowly, as expected. Once it reached 62 K, we engaged a second control heater to stabilize the SWIR focal plane temperature. After cooling stabilized, we observed that the control heaters maintained these temperatures with 47 and 234 mW dissipation for the MWIR and SWIR focal planes, respectively. Without heater control, our thermal model predicts that the telescope and MWIR focal plane would reach 56.5 - 58.4 K and 38.6 - 39.7 K respectively.
@@ 5.5 Gain and Dark Current
We produced preliminary in-flight estimates of the flat-field gain and dark current using an ensemble of images taken over varying Zodiacal brightness levels (see §3.4.3). Both estimates closely resemble measurements obtained during laboratory testing of the instrument and focal planes. As we refine the algorithms and include more data, akeson25 will contain the most recent estimates.
@@ 5.8 Stray Light Response
We are in the process of quantifying a range of stray light phenomena from in-flight observations (Dowell25). Although there are a few exceptions, the effects and levels are generally in accordance with a prelaunch design study of the stray light properties. Within the optical field of view, we can readily characterize the extended PSF to enable PSF subtraction and source masking. We observe scattered light from bright stars in narrow regions close to the frame edges, and ghost images caused by reflections in the beam splitter that are most noticeable near the 2.4 m transition.
@@ 5.9.1 Atmospheric Emission
We observe bright He line emission (Kulkarni25; Brammer14) at 1.083 m in both bands 1 and 2. The amplitude varies over the orbit and with time, ranging from a few to 50 times brighter than the Zodiacal light level. Photon noise from the line unfortunately reduces sensitivity in this band (see Figure 19). We detect low levels of spectral leakage through the LVFs from the He line, most evident at 0.84, 1.25 and 1.49 m. Other atmospheric lines are much fainter than He but detectable, most notably the atomic oxygen lines at 0.78, 0.84 and 1.13 m.
@@ 5.9.2 Shuttle Glow
Shuttle glow arises from the spacecraft interacting with the upper atmosphere. Atomic oxygen striking the leading surfaces in the direction of travel produces a glow in the optical and infrared (Viereck91) due to excited (Murad98). Emission from other molecules have been proposed in the literature, including and , which depend on the composition of the materials on the spacecraft that chemically react with atomic oxygen. Holtzclaw94 measured the emission spectrum with m resolution of the Z306 black paint used on the SPHEREx telescope, exposed to a fluence of atomic oxygen. The authors observe emission peaks at 2.8 m and 4.4 m, similar to the spectrum in Figure 21, which they attribute to vibrational emission from and . As noted in §4.2.1, we observed that shuttle glow increased strongly when the observatory was tipped forward along the direction of orbital motion. By restricting the range of zenith angles in the ram direction, we have reduced shuttle glow so that it is generally fainter than the Zodiacal sky, though a residual brightness remains.
@@ 6 Conclusions
Part of the research was carried out at the Jet Propulsion Laboratory and the California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). Work at Argonne National Laboratory was supported by the U.S. Department of Energy, Office of High Energy Physics. Argonne, a U.S. Department of Energy Office of Science Laboratory, is operated by UChicago Argonne LLC under contract no. DE-AC02-06CH11357.
The project thanks the many people who helped develop and implement hardware for the mission, including but not limited to staff at NASA, JPL, BAE, IPAC, IRSA, KASI, SpaceX, VSFB, Caltech, Viavi, Applied Aerospace Structures, UC Berkeley SSL, and Teledyne.
8 Appendix
We calculate the minimum pixel size that maintains point source sensitivity with the constraints that the number of pixels and the observing time are fixed. The sensitivity to a point source is given by
where is the effective number of pixels used to measure a point source, is the number of independent observations, is the resolving power, is Planck’s constant, is the mirror area, is the end-to-end optical efficiency, is the surface brightness of the sky set by Zodiacal light, is the solid angle of a pixel, is the integration time, is the read noise in a correlated double sample, and is the sampling interval in an SUR integration. The first term in this equation comes from photon noise, and the second from read noise. We note that the read noise term is generally an underestimate, as it only assumes white noise behavior, but additional terms contribute and become important as the integration time becomes longer. To be limited by photon noise, the first term must dominate over the second, ideally with additional margin.
The number of effective pixels in optimal photometry is given by
where is the fraction of flux that falls into pixel j. In the limit where pixels are larger than the size of the point spread function (PSF), is typically 2-3 on average, depending on the sub-pixel location of the source for each object. To maximize sensitivity, one should make the aperture A as large as possible and the spectral resolving power R as low as possible. However, the minimum resolving power was already chosen in §3.1 for sufficient redshift accuracy. Therefore, the remaining free parameters are aperture and pixel size.
SPHEREx was designed to map the entire sky with a fixed number of detector pixels, which is a non-standard optimization problem in astronomy. In this regime, the integration time and the photocurrent both scale with the pixel size as
and
where and are the photocurrent and integration time obtained with the choice of 62 pixels, and is the mirror area chosen in our design. Substituting these into the first equation, and ignoring small changes in with pixel size, we find
where and are the uncertainties in the detector current from photon noise and read noise, respectively, for our default choices of mirror area and 62 pixels. In order to be limited by photon noise, the first term in the numerator must dominate, and for our choice of parameters in bands 1-4, and the performance of the detectors, . This result means SPHEREx is comfortably limited by photon noise. However, given the strong dependence of the second term, photon noise will no longer dominate with even modest reductions in pixel size. One can also see in the photon noise limit
where increasing the aperture improves sensitivity, but only modestly.